# Low-Complexity Monitoring and Compensation of Transceiver IQ Imbalance by Multi-dimensional Architecture for Dual-Polarization 16 Quadrature Amplitude Modulation

### v1

@@ 2.2 Simulation
To confirm our proposed methods, a 36 Gbaud dual-polarization Nyquist 16QAM (DP-Nyquist-16QAM) signal with the system diagram displayed in Fig.6. In the transmitter, the full raised cosine filter is used to implement the pulse shape, and the roll-off factor is set to 0.5. The samples per symbol(SPS) are 4, and the pre-chirp filter block size is 264. Transmit a total of 216 symbols. The TX skew is set from -6 ps to 6 ps (from -T/4 to T/4 of 36 Gbaud). The scanning module scans in steps of T/16. We analyze the impact of polarization effects by the higher-order PMD model. The mean DGD is set to 0.5 ps/km1/2 in the DGD emulator. The CD coefficient is in the CD emulator. The transmission length is 100 km. The OSNR is defined assuming a bandwidth of 12.5 GHz and is set to 23 dB. In the receiver, the RX skew is set from -6 ps to 6 ps (from -T/4 to T/4T of 36 Gbaud). The signal is down-sampled to 2 SPS. The CV-MIMO tap numbers are set to 5 taps. The RV-DD-LMS tap number is set to 21 taps. The transceiver imbalances are digitally added in the DSP model of Rx and Tx. The transmission laser and the LO laser are operating at 1552.5 nm with a 100kHz linewidth and a 1 GHz frequency offset.
@@ 2.2.1 The Effect of pre-CD compensation on RX IQ Skew Estimation
First, the relation between the location of the CD compensation and the RX IQ skew estimation is tested in simulation. To clearly observe the relationship, the RX X pol IQ skew is set at 3.9 ps, and other values of transceiver IQ imbalance are set to zero in this part. We use the pre-chirp filter in the transmitter or the FDE technique in the receiver to compensate for CD. The RX skew estimation is shown in Fig. 7. According to 7, it can be found that if the CD compensation is in the receiver, the RX IQ skew estimation is close to 0 ps when the transmitter length is larger than 20 km. But if the CD compensation is in the transmitter, the RX IQ skew estimation is close to 3.9 ps, and the result is unaffected by the transmission length. This simulation result demonstrates our proposed theory that the location of CD compensation is related to RX skew estimation, and the RX IQ skew can be detected after pre-CD compensation.
@@ 2.2.2 The Effect of Transceiver IQ Imbalance on RX IQ Skew Estimation
Next, the effect of the transceiver IQ imbalance on the RX IQ skew estimation is simulated. The RX X pol skew is set from -6 ps to 6 ps. In order to comprehensively observe the imperfect factor influence on the RX X pol skew estimation error, we consider the RX X pol amplitude imbalance, phase imbalance, RX Y pol skew, and TX X pol amplitude imbalance, phase imbalance, TX X pol skew. These results are shown via 2-Dimensional(2D) planes or 3-Dimensional(3D) surfaces. Unless otherwise stated, the basic IQ imbalance is X pol IQ imbalance. For example, RX skew is RX X pol skew, TX amplitude imbalance is TX X pol amplitude imbalance, RX skew estimation error is RX X pol skew estimation error, and so on. The Tx and RX amplitude imbalance varies from -0.35 to 0.35. The TX and RX phase imbalance varies from -10 degrees to 10 degrees. The TX skew and RX Y pol skew varies from -6 ps to 6 ps, and the RX skew changes from -6 ps to 6 ps. Based on these conditions, the RX skew estimation error is shown in Fig.8. The Fig.8(a) displays the relation between RX skew estimation and RX Y pol skew. As the RX Y pol skew changes -6 ps to 6 ps, the RX skew estimation error is always kept less than 0.3 ps. This means that the RX skew estimation is unaffected by RX Y pol skew. The Fig.8(b) displays the relation between RX skew estimation and TX skew. Due to the frequency offset isolate TX skew and RX skew[7], as the TX skew changes -6 ps to 6 ps, the RX skew estimation error is always kept less than 0.3 ps. This means that the RX skew estimation is unaffected by TX skew. The reason is that the frequency offset and phase noise isolate the TX imbalance. The Fig.8(c) displays the relation between RX skew estimation, RX amplitude imbalance, and RX phase imbalance. According to Fig.8(c), the RX skew estimation error is kept less than 0.3 ps when the amplitude imbalance changes from -0.35 to 0.35, and the phase imbalance varies from -10 degrees to 10 degrees. The Fig.8(d) displays the relation between RX skew estimation, TX amplitude imbalance, and TX phase imbalance. According to Fig.8(d), the RX skew estimation error is kept less than 0.5 ps when the amplitude imbalance changes from -0.35 to 0.35, and the phase imbalance varies from -10 degrees to 10 degrees. Therefore, according to Fig.8(a) - Fig.8(d), it can be concluded that the RX skew estimation is unaffected by transceiver IQ imbalance.
2.2.3 The Effect of Transceiver IQ Imbalance and OSNR on Rough TX IQ Skew Estimation
@@
The rough TX IQ skew estimation is via error output by CV-MIMO. However, the error is affected by many factors, such as IQ imbalance, OSNR, and so on. We mainly consider the effect of transceiver IQ imbalance and OSNR on rough TX IQ skew estimation. The OSNR is set to 13 dB, 18 dB, and 23 dB, respectively. The TX skew is fixed as 2.6 ps (1.5T/16 of 36 GBaud), and the scanning module scans the TX delay at T/16 intervals. The Tx and RX amplitude imbalance varies from -0.35 to 0.35. The TX and RX phase imbalance varies from -10 degrees to 10 degrees. The RX skew is set to -6 ps, 0 ps, and 6 ps, respectively. The TX Y pol skew varies from -6 ps to 6 ps. Fig.9 displays the relation between OSNR and rough TX IQ skew estimation. First, we can find that when the OSNR is fixed, the M(e(n)) is unaffected by RX skew. Second, at OSNRs of 18 dB and 23 dB, M(e(n)) is minimum at 1.74 ps (T/16 of 36 GBaud) or 3.47 ps (2T/16 of 36 GBaud). From the scan control rule, the rough TX skew estimation is 1.74 ps or 3.47 ps. This means the residual TX skew is less than T/16 of 36 GBaud after rough TX skew compensation. However, if the OSNR is lower, such as 13 dB, the M(e(n)) can not display the correct rough TX skew estimation. The reason is that the noise is the main influence on M(e(n)) instead of TX skew.
@@
Fig.10 displays the effect of transceiver IQ imbalance on Rough TX IQ Skew Estimation. Fig.10(a) displays the relation between TX amplitude and phase imbalance and rough TX skew estimation error. Fig.10(b) displays the relation between RX amplitude and phase imbalance and rough TX skew estimation error. Fig.10(a) and Fig.10(b) display that the estimation error is kept less than 1 ps when the TX and RX phase and amplitude imbalance are changed. Fig.10(c) displays the relation between TX Y pol skew and rough TX Skew estimation error. Fig.10(c) displays that the estimation error is kept less than 1 ps when TX Y pol skew changes from -6 ps to 6 ps. According to Fig.10, it is concluded that the rough TX skew estimation error is unaffected by transceiver IQ imbalance, but is affected by the lower OSNR.
@@
It should be noted that transmitted X pol can be received as X or Y pol and vice versa. Therefore, does not allow simultaneous scanning delay on X pol and Y pol. The reason is that the scan controller is difficult to distinguish which is the X pol M(e(n)) and which is the Y pol M(e(n)), because the X pol M(e(n)) and Y pol M(e(n)) change simultaneously.
2.2.4 Performance with Transmitter IQ Imbalance
@@
In this part, we assess the performance of our proposed scheme under the transmitter IQ imbalance condition. Firstly, we assess the performance of C-R AEQ in transmitter IQ imbalance. The Tx amplitude imbalance varies from -0.35 to 0.35. The TX phase imbalance varies from -10 degrees to 10 degrees. The TX skew and TX Y pol skew vary from -6 ps to 6 ps. The Q-factor is used to display the signal quality after C-R AEQ. In order to clearly observe which is the major factor influencing the Q-factor, we study not only the 3D surfaces but also their 2D projections. Fig.11(a) displays the TX skew tolerance of C-R AEQ, and Fig.11(b) is the 2D projections of Fig.11(a). According to Fig.11(a), the maximum Q is 9.5 dB when TX skew and TX Y pol skew are 0 ps. In addition, as the TX skew approaches 6 ps when the TX Y pol skew approaches 6 ps, the Q-factor is reduced to 6.3. This means that the TX skew is one of the major influences on the Q-factor. According to Fig.11(b), the Q penalty is 0.15 dB when the Tx skew and TX Y pol skew are less than 2 ps(1/16 of symbol period). From Fig.9 and Fig.10, the scanning module makes the residual TX skew less than 1 ps. Therefore, the C-R AEQ could compensate for the residual TX skew(fine TX skew compensation) after rough TX skew compensation. Fig.11(c) shows the TX amplitude and phase imbalance tolerance of C-R AEQ, and Fig.11(d) is the 2D projections of Fig.11(c). Seen from Fig.11(c), compare to Fig.11(a), TX amplitude and phase imbalance influence on Q-factor is lower than TX skew. The Q-factor is 0.84 when the TX amplitude approaches 0.35, and the TX phase approaches 10 degrees. As shown in Fig.11(d), when the TX phase imbalance approaches 10 degrees, and there is no TX amplitude imbalance, the Q-factor remains 9.5 dB. In addition, when the TX amplitude imbalance changes from 0 to 0.35, the Q-factor changes from 9.5 to 8.4 dB. This means that the TX phase imbalance has the least influence on Q-factor after C-R AEQ, whereas the amplitude imbalance has a greater influence on Q-factor than the phase imbalance.
2.2.5 Performance with Receiver IQ Imbalance
@@
In this part, we assess the performance of our proposed scheme under the receiver IQ imbalance condition. The Rx amplitude imbalance varies from -0.35 to 0.35. The RX phase imbalance varies from -10 degrees to 10 degrees. The RX skew and RX Y pol skew vary from -6 ps to 6 ps.Fig.12(a) and Fig.12(b) display the Q-factor performance with RX skew, Y pol skew, RX phase, and amplitude imbalance. From Fig.12(a) and Fig.12(b), the maximum Q-factor is greater than 9.65 dB, and the minimum Q-factor is greater than 9.25 dB. Therefore, due to the GSOP and RX skew monitor, the Q-factor is unaffected by the RX IQ imbalance.
2.2.6 Performance with Transceiver IQ Imbalance
@@
In real conditions, the IQ imbalance is present in both the transmitter and the receiver. In addition, we want to study performance under different OSNR conditions. To simulate that condition, the TX X/Y pol skew is set to 1 ps, since the scanning module makes the TX skew less than 1 ps. The TX X/Y pol amplitude imbalance is set to 0.1. The TX X/Y pol phase imbalance is set to 5 degrees. The RX X/Y pol skew is set to 3 ps. The RX X/Y pol amplitude imbalance is set to 0.1. The RX X/Y pol phase imbalance is set to 5 degrees. These imbalanced values are from reference [17] and our experimental results. The OSNR changes from 17 dB to 23 dB. For performance comparison, we use the perfectly balanced transceiver as the reference. To clearly observe the performance of our proposed methods, the 22, 44 MIMO with 33 taps per filter and our proposed methods are used to equalize the signal, meanwhile. The result is shown in Fig.13, the Q-factor changes from 4.39 to 9.5 dB as the OSNR varies from 17 to 23 dB when there are no imbalances. The Q-factor changes from 4 to 9.2 dB as the OSNR varies from 17 to 23 dB via our proposed methods. The Q penalty is 0.3 dB when the OSNR changes from 19 dB to 23 dB. For comparison, the 22,44 MIMO Q-factor changes from 3.1 to 7.6 dB and from 3 to 8.1 dB, respectively. The Q penalty is 1.4 dB when the 44 MIMO is used at an OSNR of 23 dB. This means that our proposed methods can reduce the Q penalty caused by transceiver IQ imbalance. Comparing our proposed methods with the 44 MIMO structure, we observe that the Q factor rises by more than 1 dB at an OSNR of 23 dB.
@@ 2.2.7 Performance with DGD after C-R AEQ
Since the butterfly structure has only 5 taps per filter in C-R AEQ, it is necessary to study the performance of PMD compensation. We set the polarization rotation angle to 45 degrees, and the DGD changes from 0 to 9 ps. The DGD values refer to reference [23]. Any IQ imbalance is set to zero in this part. For comparison, the 44 MIMO with 33 taps per filter and the proposed methods are used to equalize the signal, meanwhile. The result is shown in Fig.14. It can be found that the proposed method is unaffected by DGD values ranging from 0 ps to 9 ps. In addition, the Q-factor after C-R AEQ is greater than the Q-factor after 44 MIMO by about 0.3 dB in the same DGD value. We think the reason is that the C-R AEQ further reduces the phase noise effect after CPR.

### v2

@@ 2.2 Simulation
To confirm our proposed architecture, a 36 Gbaud dual-polarization Nyquist 16QAM (DP-Nyquist-16QAM) signal with the system diagram displayed in Fig.(7). In the transmitter, the full raised cosine filter is used to implement the pulse shape, and the roll-off factor is set to 0.5. The samples per symbol(SPS) are 4, and the pre-chirp filter block size is 264. Transmit a total of 216 symbols. The TX skew is set from -6 ps to 6 ps (from -T/4 to T/4 of 36 Gbaud). We analyze the impact of polarization effects using the higher-order PMD model. The mean DGD is set to 0.5 ps/km1/2 in the DGD emulator. The CD coefficient is in the CD emulator. The transmission length is 100 km. The OSNR is defined assuming a bandwidth of 12.5 GHz and is set to 23 dB. In the receiver, the RX skew is set from -6 ps to 6 ps. The signal is down-sampled to 2 SPS. The CV-MIMO tap numbers are set to 6 taps. The TMMA-RV-MIMO tap number is set to 22 taps. The transceiver imbalances are digitally added in the DSP model of Rx and Tx. The transmission laser and the LO laser are operating at 1552.5 nm with a 100 kHz linewidth and a 1 GHz frequency offset.
@@ 2.2.1 The Effect of pre-CD compensation on RX IQ Skew Estimation
First, the relation between the CD compensation and the RX IQ skew estimation is tested in simulation. In Fig. 8, the estimation results are plotted as a function of fiber length. To clearly observe the relationship, the RX X pol IQ skew is set at 3.9 ps, and other values of transceiver IQ imbalance are set to zero in this part. We use the pre-chirp filter in the transmitter or the FDE technique in the receiver to compensate for CD. It should be noted that the transmission length is less than 10 km, the CD compensation is not performed. The RX skew estimation is shown in Fig. 8. According to Fig. 8, it can be found that if the CD compensation is in the receiver, the RX IQ skew estimation is close to 0 ps when the transmitter length is larger than 20 km. But if the CD compensation is in the transmitter, the RX IQ skew estimation is close to 3.9 ps, and the result is unaffected by the transmission length. In general, receiver IQ skew is uncorrelated with the transmission length. However, from Fig. 8, if the optical signal has a larger CD and the CD is compensated in the receiver, the receiver IQ skew cannot be detected by GSMA. This simulation result demonstrates our proposed theory that the location of CD compensation is related to RX skew estimation, and the RX IQ skew can be detected after pre-CD compensation.
@@ 2.2.2 The Effect of Transceiver IQ Imbalance on RX IQ Skew Estimation
Next, the effect of the transceiver IQ imbalance on the RX IQ skew estimation is simulated. The RX X pol skew is set from -6 ps to 6 ps. In order to comprehensively observe the imperfect factor influence on the RX X pol skew estimation error, we consider the RX X pol amplitude imbalance, phase imbalance, RX Y pol skew, and TX X pol amplitude imbalance, phase imbalance, TX X pol skew. These results are shown via 2-Dimensional(2D) planes or 3-Dimensional(3D) surfaces. Unless otherwise stated, the basic IQ imbalance is X pol IQ imbalance. For example, RX skew is RX X pol skew, TX amplitude imbalance is TX X pol amplitude imbalance, RX skew estimation error is RX X pol skew estimation error, and so on. The Tx and RX amplitude imbalance varies from -0.35 to 0.35. The TX and RX phase imbalance varies from -10 degrees to 10 degrees. The TX skew and RX Y pol skew varies from -6 ps to 6 ps, and the RX skew changes from -6 ps to 6 ps. Based on these conditions, the RX skew estimation error is shown in Fig.9. The Fig.9(a) displays the relation between RX skew estimation and RX Y pol skew. As the RX Y pol skew changes -6 ps to 6 ps, the RX skew estimation error is always kept less than 0.25 ps. This means that the RX skew estimation is unaffected by RX Y pol skew. The Fig.9(b) displays the relation between RX skew estimation and TX skew. Due to the frequency offset isolate TX skew and RX skew[7], as the TX skew changes -6 ps to 6 ps, the RX skew estimation error is always kept less than 0.3 ps. This means that the RX skew estimation is unaffected by TX skew. The reason is that the frequency offset and phase noise isolate the TX imbalance. The Fig.9(c) displays the relation between RX skew estimation, RX amplitude imbalance, and RX phase imbalance. According to Fig.9(c), the RX skew estimation error is kept less than 0.25 ps when the amplitude imbalance changes from -0.35 to 0.35, and the phase imbalance varies from -10 degrees to 10 degrees. The Fig.9(d) displays the relation between RX skew estimation, TX amplitude imbalance, and TX phase imbalance. According to Fig.9(d), the RX skew estimation error is kept less than 0.5 ps when the amplitude imbalance changes from -0.35 to 0.35, and the phase imbalance varies from -10 degrees to 10 degrees. Therefore, according to Fig.9(a) - Fig.9(d), it can be concluded that the RX skew estimation is unaffected by transceiver IQ imbalance.
2.2.3 The Effect of Transceiver IQ Imbalance and OSNR on TX IQ Skew Estimation
@@
The TX IQ skew estimation is via the output by CV-MIMO. However, the error is affected by many factors. We mainly consider the effect of transceiver IQ imbalance and OSNR on TX IQ skew estimation. The OSNR is set to 13 dB, 18 dB, and 23 dB, respectively. The TX skew is fixed as 2.6 ps (1.5T/16 of 36 GBaud), and SDA scans the TX delay at T/16 intervals. The Tx and RX amplitude imbalance varies from -0.35 to 0.35. The TX and RX phase imbalance varies from -10 to 10 degrees. The RX skew is set to -6 ps, 0 ps, respectively. The TX Y pol skew varies from -6 to 6 ps. Fig.10 displays the relation between OSNR and TX IQ skew estimation. At OSNRs of 18 dB and 23 dB, is minimum at 1.74 ps (T/16 of 36 GBaud) or 3.47 ps (2T/16 of 36 GBaud). From the scan control rule, the TX skew estimation is 1.74 ps or 3.47 ps. However, at the OSNR of 13 dB, the can not display the correct TX skew estimation. The reason is that the noise is the main influence on instead of TX skew.
@@
Fig.11 displays the effect of transceiver IQ imbalance on TX IQ skew estimation. Fig.11(a) and Fig.11(b) display that the estimation error is kept less than 0.9 ps when the TX and RX phase and amplitude imbalance are changed. Fig.11(c) displays that the estimation error is kept less than 0.9 ps when TX Y pol skew changes from -6 ps to 6 ps. According to Fig.11, it is concluded that the TX skew estimation error is unaffected by transceiver IQ imbalance.
@@
It should be noted that transmitted X pol can be received as X or Y pol and vice versa. Therefore, simultaneous scan delay on both X and Y polarizations is not allowed.
2.2.4 Performance with Transmitter IQ Imbalance after C-R AEQ
@@
In this part, we assess the performance of C-R AEQ under the transmitter IQ imbalance condition. The Tx amplitude imbalance varies from -0.35 to 0.35. The TX phase imbalance varies from -10 degrees to 10 degrees. The TX skew and TX Y pol skew vary from -6 ps to 6 ps. The Q-factor is used to display the signal quality after C-R AEQ. In order to clearly observe which is the major factor influencing the Q-factor, we study not only the 3D surfaces but also their 2D projections. Fig.12(a) displays the TX skew tolerance of C-R AEQ, and Fig.12(b) is the 2D projections of Fig.12(a). According to Fig.12(a), the maximum Q is 9.5 dB when TX skew and TX Y pol skew are 0 ps. As the TX skew approaches 6 ps when the TX Y pol skew approaches 6 ps, the Q-factor is reduced to 6.3. This means that the TX skew is one of the major influences on the Q-factor. According to Fig.12(b), the Q penalty is 0.15 dB when the Tx skew and TX Y pol skew are less than 2 ps(1/16 of symbol period). From Fig.10 and Fig.11, the SDA makes the residual TX skew less than 1 ps. Fig.12(c) shows the TX amplitude and phase imbalance tolerance of C-R AEQ, and Fig.12(d) is the 2D projections of Fig.12(c). Seen from Fig.12(c), compare to Fig.12(a), TX amplitude and phase imbalance influence on Q-factor is lower than TX skew. The Q-factor is 8.4 when the TX amplitude approaches 0.35, and the TX phase approaches 10 degrees. As shown in Fig.12(d), when the TX phase imbalance approaches 10 degrees, and there is no TX amplitude imbalance, the Q-factor remains 9.5 dB. In addition, when the TX amplitude imbalance changes from 0 to 0.35, the Q-factor changes from 9.5 to 8.4 dB. This means that the TX phase imbalance has the least influence on Q-factor after C-R AEQ, whereas the amplitude imbalance has a greater influence on Q-factor than the phase imbalance.
2.2.5 Performance with Receiver IQ Imbalance after C-R AEQ
@@
In this part, we assess the performance of C-R AEQ under the receiver IQ imbalance condition. The Rx amplitude imbalance varies from -0.35 to 0.35. The RX phase imbalance varies from -10 degrees to 10 degrees. The RX skew and RX Y pol skew vary from -6 ps to 6 ps.Fig.13(a) and Fig.13(b) display the Q-factor performance with RX skew, Y pol skew, RX phase, and amplitude imbalance. From Fig.13(a) and Fig.13(b), the maximum Q-factor is greater than 9.65 dB, and the minimum Q-factor is greater than 9.25 dB. Therefore, due to the GSOP and GSMA, the Q-factor is unaffected by the RX IQ imbalance.
2.2.6 Performance with Transceiver IQ Imbalance after C-R AEQ
@@
In real conditions, the IQ imbalance is present in both the transmitter and the receiver. Moreover, we aim to study performance under different OSNR conditions. To simulate that condition, the TX X/Y pol skew is set to 1 ps, since the SDA makes the TX skew less than 1 ps. The TX X/Y pol amplitude imbalance is set to 0.1. The TX X/Y pol phase imbalance is set to 5 degrees. The RX X/Y pol skew is set to 0.25 ps, since the GSMA makes the RX skew less than 0.25 ps. The RX X/Y pol amplitude imbalance is set to 0.1. The RX X/Y pol phase imbalance is set to 5 degrees. These imbalanced values are from our experimental results. The OSNR changes from 17 dB to 23 dB. The equalization performance of our proposed methods is compared against conventional 2×2 and 4×4 MIMO (33 taps per filter), using a perfectly balanced transceiver as the reference. Results are shown in Fig. 14. First, under balanced conditions, the Q-factor varies from 4.39 to 9.5 dB as the OSNR varies from 17 to 23 dB. Second, under imbalanced conditions, the Q-factor changes from 4 to 9.2 dB as the OSNR varies from 17 to 23 dB via our proposed methods. The Q penalty is 0.3 dB when the OSNR changes from 19 dB to 23 dB. For comparison, the 22 and 44 MIMO Q-factor changes from 3.1 to 7.6 dB and from 3 to 8.1 dB, respectively. The Q penalty is 1.4 dB when the 44 MIMO is used at an OSNR of 23 dB. This means that our proposed methods can reduce the Q penalty caused by transceiver IQ imbalance. Comparing our proposed methods with the 44 MIMO structure, we observe that the Q factor rises by more than 1 dB at an OSNR of 23 dB.
@@ 2.2.7 Performance with DGD after C-R AEQ
Since the CV-MIMO only has 6 taps per filter, it is necessary to study the performance of PMD compensation. We set the DGD to vary from 0 to 9 ps. All IQ imbalance is set to zero in this part. For comparison, the 44 MIMO with 33 taps per filter and the C-R AEQ are used to equalize the signal. As shown in Fig.15, the C-R AEQ is unaffected by DGD values ranging from 0 ps to 9 ps. Moreover, the Q-factor after C-R AEQ is greater than the Q-factor after 44 MIMO by about 0.3 dB in the same DGD value.